the lengh of a rectangle is 10 feet longer than the with. if the perimeter is 96 feet find the lengh and the with of the perimeter
Answer by addingup(3677)

So you say that:
L = W+10
The perimeter is:
2L + 2W = 96 And since we just said that L = W+10, let's substitute like this:
2(W+10) + 2W = 96 Multiply of the left:
2W + 20 + 2W = 96 Add W's on left and subtract 20 on both sides:
4W = 76 Divide both sides by 4:
W = 19 <-- This is the Width. And we said the length is 10 more:
L = 19+10 = 29
Proof:
2L + 2W = 96; 2(29) + 2(19) = 96; 58 + 38 = 96; 96 = 96 We've got the right answer.